OpenStudy (swaugh):

Which equation is not equivalent to the others? zero point four x equals two point four three x equals one point eight zero point five x equals three x equals six

OpenStudy (calculusxy):

A. \[0.4x = 2.4\] B. \[3x = 1.8\] C. \[0.5x = 3\] D. \[x = 6\] Solve for x by dividing on options a, b, and c.

OpenStudy (swaugh):

but i dont know what x equals

OpenStudy (calculusxy):

for example, we will use option a. 0.4x = 2.4 to solve for x we need to use the division property of equality and do 2.4/0.4 to get x.

OpenStudy (calculusxy):

can you do that see what x equals to?

OpenStudy (calculusxy):

\[\large 2.4 \div 0.4 = x \]

OpenStudy (swaugh):

is it b

OpenStudy (swaugh):

b = 1.66666667

OpenStudy (calculusxy):

Yes it is b but you have it for the wrong reason. for option a, x = 6 (2.4/0.4 = 6) b, x = 0.6 (1.8/3 = 0.6) c, x = 6 (3 / 0.5 = 6) d, x = 6

OpenStudy (calculusxy):

sorry i meant 0.6

OpenStudy (calculusxy):

while the others have simply 6

OpenStudy (swaugh):

the answer is b

OpenStudy (calculusxy):

Yes

OpenStudy (swaugh):

thank you

OpenStudy (calculusxy):

you're welcome